OpenStudy (stranger.tz):

A tank can be filled by tap A in 2 hours and by tap B in 6 hours .how long will it take both tape to fill the tank?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{A=\dfrac{1}{t_1}\\~\\ A=\dfrac{1}{2}\\~\\ B=\dfrac{1}{t_2}\\~\\ B=\dfrac{1}{6}\\~\\ A+B=\dfrac{1}{2}+\dfrac{1}{6}\\~\\ =\dfrac{3}{6}+\dfrac{1}{6}\\~\\ =\dfrac{4}{6}\\~\\ =\dfrac{2}{3}\\~\\ =\dfrac{1}{\dfrac{3}{2}}\\~\\ =\dfrac{1}{1.5}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

so \(\Large A+B\) will use \(\Large 1.5~~hrs\)

OpenStudy (stranger.tz):

Thanks